T S 35 25 25 S 15 For the independence model, the expected frequencies are given by 30 20 720 25 30 25 15 30 25 25 2525 30 30 20 20
Answer #1

The Expected frequencies are  

Obs Freq. Exp. Freq.
T T' Total T T' Total
S 35 25 60 =(50*60)/100 =30 30 60
S' 15 25 40 =(50*40)/100 =20 20 40
Total 50 50 100 50 50 100

Answer: Option

30 30